How high a column of air would be necessary to cause a barometer to read 76 cm of hg if the atmosphere were of uniform density 1.2 kg/m³? The density of hg= 13.6*kg/m³

Grade:11

2 Answers

Vikas TU
6 years ago
Dear Student,
density of hg= mass/ (area*76);
Density of air= mass/(area*x)
Dividing 1 with 2,
13.6/1.2=x/76;
x= 861.33cm.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)
Aman
4 years ago
Given: dair=1.2 kg/m^3 
dHg=13.6×10^3
 
Let height of air column = Hair
And height of Hg =Hhg =76cm of Hg =0.76m of Hg
 
Now 
We know that 
               dair×Hair×g = dHg×Hhg×g
 
                 1.2×Hair =13.6×10^3 × 0.76
                         Hair = 8613m